OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 55 mi / h = ____ ft / s

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

not a clue

OpenStudy (skullpatrol):

$$\Huge 55\frac{mi}{h}----> \frac{ft}{s}$$

OpenStudy (skullpatrol):

Convert miles into feet in the numerator and hours into seconds in the denominator, right?

OpenStudy (anonymous):

so how would i do that ?

OpenStudy (anonymous):

ohhhh wait would it be 80.7 ?

OpenStudy (skullpatrol):

How many seconds make up an hour?

OpenStudy (anonymous):

3600 seconds?

OpenStudy (skullpatrol):

Correct! :-)

OpenStudy (anonymous):

:D

OpenStudy (skullpatrol):

$$\Huge 55\frac{mi}{{h}\frac{3600s}{h}}----> \frac{ft}{s}$$

OpenStudy (anonymous):

806.7 ^-^

OpenStudy (skullpatrol):

How many feet make up a mile?

OpenStudy (anonymous):

5280

OpenStudy (skullpatrol):

Greeeeat :-)

OpenStudy (anonymous):

:DDDDDDDDDDDDDDDDDD

OpenStudy (skullpatrol):

$$\Huge 55\frac{mi\frac{5280ft}{mi}}{{h}\frac{3600s}{h}}----> \frac{ft}{s}$$